# Object Re-identification via Spatial-temporal Fusion Networks and Causal Identity Matching

### v1

@@ VII-A Settings and evaluation metrics
Unlike the appearance-centric (1-to-all) evaluation settings in Section. VI, in this experiment, we perform ID-to-ID evaluation to consider the real-world scenario. This means that the ReID performances should be evaluated only once for each identity between cameras. For example, each identity has a path across cameras, e.g., $c_{1}\rightarrow c_{2}\rightarrow c_{3}\rightarrow c_{4}$; thus, ReID methods should correctly track the path based on the ReID between adjacent cameras. For ReID between cameras in the real world, only the rank-1 matching result is valid. Therefore, we considered the rank-1 result with a similarity value over $\theta_{c}$ as a positive matching result. We empirically set $\theta_{c}=0.6$.
To evaluate the performances, the mean Average Precision (mAP) and $F_{1}$ score of predicted object paths are measured. Vehicle-3I includes 1,630 training IDs and 408 test IDs, the same as previous experiments in Section. VI. We evaluate our methods in real-world scenarios using Vehicle-3I dataset. For evaluation, each test query progressively performs ReID according to the estimated connection matrix. ReID for each query continues until no identities are satisfying $\theta_{c}>0.6$ between the test query and its gallery. In this way, the ReID of all test queries is performed, and the performances (mAP, $F_{1}$) are measured.
Figure 12: Appearances of the target object. The appearances in camera 7 (C7) are the initial query. Blue box denotes appearances in difference cameras (C6,C1,C3) after performing Causal Identity Matching (CIM).
@@ VII-B Effects of Causal Identity Matching (CIM)
The connection matrix for Vehicle-3I dataset can be estimated using the paths of training identities. The size of connection matrix $\mathbf{C}_{3I}$ for Vehicle-3I is $11\times 1$ 11. This connection matrix is quite sparse, with only a few adjacent connections between cameras. This means that when performing ReID for the target object, the entire distributions $p_{ij}\left(t\right)$ are not needed; only some adjacent $p_{ij}\left(t\right)$ are considered at each ReID step. Furthermore, as shown in Fig. 12, the appearances of the target object become much richer through CIM, allowing for more accurate ReID. Although a single camera can only capture fixed poses of the object, progressive query merging easily overcomes this limitation. CIM increases the overall efficiency and accuracy of the ReID system, as summarized in Table VI. This is a very natural procedure, but note that most recent studies have not closely examined such real-world conditions and causality for ReID.
@@
TABLE VI: ReID performances without and with CIM.
@@
As summarized in Table. VII, we compared several top-$k$ multi-shot matching methods including $k=5,10,20$, and tested the ReID models FastReID [5] and ours.
In both methods (FastReID [5], Ours), the single-shot matching strategy performed lowest accuracies due to limited appearance cue. Furthermore, using a single appearance for each identity does not correspond to real-world scenarios. Compared to the ‘single-shot’, multi-shot (min) and multi-shot (avg.) strategies achieved higher performances. This results support that using multiple appearance improves the ReID performances. However, multi-shot (max) can be highly biased in the best similarity score between two sets of appearances. Meanwhile, multi-shot (avg.) can consider multiple comparisons, but the final score can be smoothed out or polluted by unreliable appearances.
On the other hand, the proposed Top-$k$ matching is neither too biased nor too smoothed, and it produces a stable score even when the number of compared appearances changes by appearance management as in proposed in Section IV-B. Regardless of the $k$ value, the proposed Top-$k$ multi-shot matching methods achieved higher performance than other strategies such as single and multi-shot (max, avg.) in both methods (FastReID [5], Ours). It performed 94.95% mAP and 95.19 % $F_{1}$ score when the $k=5$.
@@ VII-D Effects of appearance management
The complexity of multi-shot matching between two appearance sets is $O\left(N_{m}N_{n}\right)$, where $N_{m},N_{n}$ are the appearance numbers of $m$- and $n$-th objects. This roughly follows a complexity of $O\left(N^{2}\right)$ and needs to be improved for efficiency. Furthermore, among the multiple appearances, many duplicated and polluted appearances lead to performance degradation, as described in Section IV-B. Based on the proposed appearance unreliability score $\mathcal{U}\left(\mathbf{I}^{m}\right)$, we adjusted the number of appearances in each object.
In Table VIII, we tested different settings of the proposed appearance management and performed Top-$5$ multi-shot matching. The App. selection ratio means the proportion of collected appearances used for ReID. For example, we can use only a few reliable appearances such as $5-50\%$ according to the unreliability score $\mathcal{U}\left(\mathbf{I}^{m}\right)$. App. selection ratio of ‘ALL’ used all collected appearances without performing the proposed appearance management. The proposed appearance management improved the ReID performances about 2% and 1.32% mAP scores for each method (FasetReID, Ours), when they only use $5\%$ reliable appearances, respectively. These results indicate that, regardless of the ReID methods used, ReID performance can be improved with only a few selected appearances. In addition, the matching complexities are dramatically reduced. Additionally, matching complexities are significantly reduced. For example, when selecting only 5% of appearances, 99.75% of redundant calculations can be eliminated, while maintaining or even enhancing performance.
@@
TABLE VII: Performance comparisons of various matching strategies on the Vehicle-3I dataset. Single-shot and multi-shot (max, avg.) are the traditional ReID matching strategies, and multi-shot (top-$k$) is our proposed matching method. The best performances are marked in bold.
@@
TABLE VIII: ReID performance according to the App. selection ratio for the proposed appearance management on the Vehicle-3I dataset. Best performances are marked in bold.
@@ VIII Conclusion
In this study, we proposed a ReID framework based on a spatial-temporal fusion network and causal identity matching (CIM), which considers the issues in real-world scenarios. The proposed framework estimates camera network topology and combines appearance and spatial-temporal similarities to reduce appearance ambiguity. To this end, we proposed an adaptive Parzen window for reliable topology estimation and FusionNet for optimal similarity aggregation. The proposed framework demonstrates outstanding performances on the VeRi776 [14], Vehicle-3I and Market-1501 [13] datasets, achieving 99.70%, 89.22%, and 99.11% rank-1 accuracy and 91.71%, 83.62%, and 95.5% mAP, respectively. The results indicate that incorporating spatial and temporal information in ReID can improve the accuracy of appearance-based methods and effectively address appearance ambiguity.
Additionally, we proposed causal identity matching (CIM) to consider the issues in real-world scenarios. The proposed CIM uses adjacent cameras and transition time distributions to dynamically assign gallery sets for target objects to include only relevant objects and progressively merge queries to improve ReID accuracy and variety. In Vehicle-3I dataset, proposed CIM achieved outstanding performances by 94.95% mAP and 95.19% F1 score. The results demonstrate that the proposed CIM is effective in real-world surveillance scenarios beyond the laboratory setting. Although the proposed ReID framework has lightweight networks, it can effectively perform ReID tasks. Furthermore, the proposed CIM is flexible enough to use other ReID frameworks and appearance-based methods.

### v2

@@ VII-A Settings and evaluation metrics
Unlike the appearance-centric (1-to-all) evaluation settings in Section. VI, in this experiment, we perform ID-to-ID evaluation to consider the real-world scenarios. This means that ReID performances should be evaluated only once for each identity between cameras. For example, each identity has a path across cameras, e.g., $c_{1}\rightarrow c_{2}\rightarrow c_{3}\rightarrow c_{4}$; thus, ReID methods should correctly track the path based on the ReID between adjacent cameras. In real-world ReID scenarios, only the rank-1 matching result is considered valid. Therefore, we regarded the rank-1 result with a similarity value over $\theta_{c}$ as a positive match. We empirically set $\theta_{c}=0.6$.
To evaluate the performances, the mean Average Precision (mAP) and $F_{1}$ score of predicted object paths are measured. The Vehicle-3I dataset includes 1,630 training IDs and 408 test IDs, consistent with the previous experiments in Section. VI. We evaluated our methods in real-world scenarios using the Vehicle-3I dataset. For the evaluation, each test query was performed progressively using ReID according to the estimated connection matrix. The ReID process for each query continued until no identities satisfied the condition $\theta_{c}>0.6$ between the test query and its gallery. In this way, the ReID of all test queries is performed, and the performances (mAP, $F_{1}$) are measured.
Figure 12: An example of the progressive query merging. The appearances in camera 7 (C7) are the initial query. The blue box denotes merged appearances in difference cameras (C6,C1,C3) after performing Causal Identity Matching (CIM).
@@ VII-B Effects of Causal Identity Matching (CIM)
The connection matrix for the Vehicle-3I dataset can be estimated using the paths of training identities. The size of the connection matrix $\mathbf{C}_{3I}$ for Vehicle-3I is $11\times 1$ 11. This matrix is sparse, with only a few adjacent connections between cameras. This means that when performing ReID for the target object, the entire distributions $p_{ij}\left(t\right)$ are not needed; only some adjacent $p_{ij}\left(t\right)$ are considered at each ReID step. Furthermore, as shown in Fig. 12, the appearances of the target object become much richer through CIM and progressive query merging, allowing for more accurate ReID. Although a single camera can capture only fixed poses of the object, the query merging stage in CIM easily overcomes this limitation. It increases the overall efficiency and accuracy of the ReID system, as in Table V. This is a very natural procedure, but note that most recent studies have not closely examined such real-world conditions and causality for ReID.
@@
TABLE V: ReID performances without and with CIM.
@@
As summarized in Table. VI, we compared several top-$k$ multi-shot matching methods including $k=5,10,20$, and tested the ReID models FastReID [5] and ours.
In both methods (FastReID [5], Ours), the single-shot matching strategy performed lowest accuracies due to the limited appearance cues. Moreover, using a single appearance for each identity does not correspond to real-world scenarios. Compared to the ‘single-shot’, multi-shot (min) and multi-shot (avg.) strategies achieved higher performances. This results support that using multiple appearances improves ReID performances. However, multi-shot (max) can be highly biased in the best similarity score between two sets of appearances. Meanwhile, multi-shot (avg.) can consider multiple comparisons, but the final score can be smoothed out or polluted by unreliable appearances.
On the other hand, the proposed Top-$k$ matching is neither too biased nor too smoothed, and it produces a stable score even when the number of compared appearances changes by appearance management as proposed in Section IV-B. Regardless of the $k$ value, the proposed Top-$k$ multi-shot matching methods achieved higher performance than other strategies such as single and multi-shot (max, avg.) in both methods (FastReID [5], Ours). It achieved a 94.95% mAP and a 95.19 % $F_{1}$ score when $k=5$.
@@ VII-D Effects of appearance management
The complexity of multi-shot matching between two appearance sets is $O\left(N_{m}N_{n}\right)$, where $N_{m},N_{n}$ are the appearance numbers of $m$- and $n$-th objects. This roughly follows a complexity of $O\left(N^{2}\right)$ and needs to be improved for efficiency. Additionally, many duplicated and noisy appearances degrade performance, as discussed in Section IV-B. To address this, we utilized the proposed appearance unreliability score $\mathcal{U}\left(\mathbf{I}^{m}\right)$ to adjust the number of appearances for each object.
In Table VII, we tested different settings for the proposed appearance management and performed Top-$5$ multi-shot matching. The App. selection ratio means the proportion of collected appearances used for ReID. For example, we can use only a few reliable appearances such as $5-50\%$ according to the unreliability score $\mathcal{U}\left(\mathbf{I}^{m}\right)$. App. selection ratio of ‘ALL’ used all collected appearances without performing the proposed appearance management. The proposed appearance management improved ReID performances by approximately 2% and 1.32% in mAP scores for the FasetReID and our method, respectively, when only $5\%$ of reliable appearances were utilized. These results indicate that, regardless of the ReID methods used, ReID performance can be improved with only a few selected appearances. In addition, the matching complexities are dramatically reduced. Additionally, matching complexities are significantly reduced. For example, when selecting only 5% of appearances, 99.75% of redundant calculations can be eliminated, while maintaining or even enhancing performance.
@@
TABLE VI: Performance comparisons of various matching strategies on the Vehicle-3I dataset. Single-shot and multi-shot (max, avg.) are the traditional ReID matching strategies, and multi-shot (top-$k$) is our proposed matching method. The best performances are marked in bold.
@@
TABLE VII: ReID performances according to the App. selection ratio for the proposed appearance management. The best performances are marked in bold.
@@ VIII Conclusion
This study proposes a ReID framework integrating a spatial-temporal fusion network with causal identity matching (CIM) to address real-world challenges. The framework estimates camera network topology and combines appearance with spatial-temporal similarities to reduce appearance ambiguity. To this end, we proposed an adaptive Parzen window for reliable topology estimation and FusionNet for optimal similarity aggregation. The proposed framework achieves remarkable performance on the VeRi776 [14], Vehicle-3I and Market-1501 [13] datasets, with rank-1 accuracies of 99.70%, 89.22%, and 99.11% and mAP scores of 91.71%, 83.62%, and 95.5% mAP, respectively. The results indicate that incorporating spatial and temporal information in ReID can improve the accuracy of appearance-based methods and effectively address appearance ambiguity.
Additionally, we proposed causal identity matching (CIM) to consider the issues in real-world scenarios. The proposed CIM leverages adjacent cameras and transition time distributions to dynamically assign gallery sets for target objects to include only relevant objects and progressively merge queries to improve ReID accuracy and variety. On the Vehicle-3I dataset, CIM achieved notable performances with a 94.95% mAP and 95.19% F1 score. The results demonstrate that the proposed CIM is effective in real-world surveillance scenarios beyond the laboratory setting. Although the proposed ReID framework utilizes lightweight networks, it effectively performs ReID tasks. The proposed CIM is flexible and can be integrated with other ReID frameworks and appearance-based methods.